### v2

@@ 3 Problem introduction and setup
The uniformly separated measurements around the cylinder can be used to determine the drag coefficient by summation of discrete measurements as approximate integration. For further details about the dimensions of the setup, refer to Figure 2.
@@
$u(y)=\frac{4U_{y}(0.41-y)}{0.41^{2}}$
$y$ is the y-coordinates, and $U_{y}$ is 1.5 in this scenario, instead of the sinusoidal profile in the test problem as provided by Turek et al.[50, 51]. Furthermore, the outflow is the rightmost wall. The upper, lower, and obstacle walls all have a non-slip condition ($u$=0) as presented in Figure 2.
4 Jets Configuration
Two jets with blowing and suction control are used to manipulate the flow. The first jet (referred to as Jet 1) is at the top of the circular base of the cylinder (at coordinate (0.2, 0.25)), and the second jet (Jet 2) is at the bottom (at coordinate (0.2, 0.15)). The width of the jets is small, at 0.25 percent of the diameter. The jets can perform blowing and suction independently, meaning blowing and suction can happen simultaneously. A reinforcement learning algorithm is applied to control the blowing and suction of the jets. More information about the execution of the simulation will be provided in the next section.
5 Feedback Formulation
@@
To measure the SPLs, pressure is recorded from the pressure field provided in the simulation with surfaces of closely located sensors around the cylinder, 0.05m away from the container’s walls (refer to Figure 3). The sensor surfaces enclose the vortex street created by the flow and therefore can give a more accurate reflection of the varying pressure field along the vortex street and it helps us determine the static pressure level, which are essential for the calculation of the SPLs. We set 2000 sensor points horizontally and 500 sensor points vertically on each side. The pressure of each point at a particular time is then extracted from the pressure field created by the simulation. The upper and lower sensor surfaces are mainly concerned as they cover the length of the vortex street, which will be the source of most of the noise generation. These vortices are born from instabilities in the bottom and top regions of flow separation alternatively. Hence, having a distinction between the top horizontal sensor array and the bottom horizontal sensor array helps in understanding the vortex periodicity. Pressure values of every sensor point at each time step are recorded and passed through a function to convert to the relative $SPL_{i}$ for each sensor and effective SPL for the system $SPL_{eff}$ using the formulae below.
$SPL_{i}=20\mathrm{log}\frac{|p_{i}-p_{avg}|_{rms}}{p_{avg}}$
$SPL_{eff}=10\mathrm{log}\frac{\sum_{i=1}^{n}|p_{i}-p_{avg}|^{2}_{rms}}{p^{2}_{avg}}$
@@ 6 Control Strategies
6.1 Stage 1: Explore the optimal range of jet values
@@
6.1.1 Set-up
It is challenging to determine what velocity value of the jet is able to influence the flow as we desired. Therefore, build-up strategy is used for the DQN to explore the value that can lower the SPLs. The velocity values of the jets are generated based on the values of the previous time step. The increments and decrements include ±0.01, ±0.05, ±0.1, ±0.5 and 0, so there shall be 81 combinations of actions for the DQN algorithm to manage. For example, if both jets initially have the values of 0.1, the algorithm will have the option to decrease Jet 1 by 0.01, 0.05, 0.1, 0.5 or keep it constant, and likewise for Jet 2. This results in a large number of actions for the DQN to explore, which can potentially hinder the ability to optimize the reward as the optimization problem needs more variables to be tuned and the gradient surface becomes rough and noisy, though the capability of optimization with more variables is better. Hence, fewer controlling variables are typically preferred. The velocity values are kept constant between the two interventions. The input of the DQN is a vector with a dimension of 1x4. The first two are for the SPLs of upper and lower sensor surfaces and the last two are the velocity values of Jet 1 and Jet 2. In this stage, we follow the model in Figure 2, which allows the simulation to pass the data (reward, action, states) to the DQN at every time step. However, the jet velocity and therefore the action are kept constant every 50 time steps.
The reward returned is determined by a function in the simulation. In this task, the simulation learns how to reduce approximately 3 - 5 dB and create a convergence in the process. A simulation runs without any jet intervention results in two oscillations. While the lower SPL peaks at roughly 74.5 dB, the upper SPL’s peak is about 0.25 dB lower. Moreover, the range of the SPL is from 71.7 to 74.5 dB (2.8 from maximum to minimum). Refer to Figure 5 for further information. Meanwhile, the drag coefficient is also measured due to its popularity in wake control, so this measurement can be used as a means of verification for the noise reduction method. The drag experienced initially is large due to the direction of flow at the beginning. When the flow stabilizes, the plot suggests it has an oscillatory behaviour, with a maximum and a minimum at approximately 3.185 and 3.123, respectively.
@@ 6.1.1 Set-up
6.1.2 Results
After applying the jet intervention and running the simulation to its completion, the result is plotted in Figure 6. From t = 4 to t = 6 on the x-axis, the SPL is stable because there is no intervention yet. However, from t = 6, there are many fluctuations in the SPL due to the exploration of the DQN algorithm. Many large values of jet values are chosen, which leads to observable discrepancies before the 8-second mark. The overall SPL in this time range is still relatively the same as when the jet is not turned on, but there is a small surge near t = 8, demonstrating the model has been able to achieve a jet pattern that can affect the SPL oscillations.
From t = 8 to t = 10, the SPL continues to rise to a peak of 75 dB, then gradually falls over time. It is also evident that there are some interventions that can bring the SPL down, preventing it from increasing drastically. The act of exploration searches for optimal policy control by letting random actions to learn the optimality like mutation steps in the genetic algorithm and simultaneously exploits the learning with the actions that reduce the SPL by preventing the random actions letting more actions from optimal policy to learn optimality. This causes the initial SPL to rise and then gradually drop[52, 53]. The SPL keeps its decreasing tendency until t = 16, as the interventions are noticeably fewer as the DQN agent slowly enters the exploiting phase.
@@ 6.1.2 Results
Looking at the velocity plot in Figure 6b, it can be seen that a velocity with a magnitude of around 1.0 can influence the flow to the extent we desire. Along with results from other trials, one of which will be presented in Figure 7, it can be deduced that jet velocity with a magnitude above 3 shall likely cause simulation failure due to the breakdown of the PDE solver at high jet velocities due to the instability of the discretization scheme because of the high courant number near the jets. Essentially, it is a computational limitation and it can be tackled by using higher-order numerical schemes or finer discretization. In Figure 7a, although the SPL is at the desired value, the simulation is stopped at just past t =10 (5000 mark on the x-axis). So, the lower value of the jet is at -3.5 and that is not ideal for the simulation. This threshold will be used to limit the jet speed in Stage 2 for the test cases.
6.2 Stage 2: Testing with definite jet velocity values
6.2.1 Set-up
@@
Based on the result of the previous section, we set a deterministic DQN algorithm with blowing and suction of three different values. In this stage, the jet velocity is no longer built up from the previous time step but rather has a distinct, fixed set of values. Two different test cases are presented: [±1.50, ±2.25, ±2.75] and [±2.00, ±2.50, ±3.00] (referred to as Test case 1 and Test case 2 respectively), to verify the result. The expectation from this stage is similar to the one before, which is lowering from 3 - 5 dB approximately. The reward calculation function is kept the same as in the previous stage (as in Table 1), but the reward calculation process is different. While the previous stage fully adopts the Markov model (Figure 1), the strategy for this stage is accumulating the rewards between two interjections and averaging them, then returning to the DQN. Lastly, the 50-divisible state is returned to the DQN. Both test cases will implement this model. Refer to Figure 8 for further details.
@@
6.2.2 Results
The resulting SPL of the upper and lower sensors are displayed in Figure 9 for both cases. Test case 1, using velocity values with lower magnitudes, converges to an SPL of just below 70 dB. Meanwhile, Test case 2 demonstrates a slightly lower SPL than that of Test case 1, and it seems to converge earlier as well. Also, the amplitude of oscillations is evident to have reduced.
On the other hand, the instantaneous drag coefficient in each case is calculated and the result is impressive(refer to Figure 10). The system experiences a large drag force at the start due to the direction of the flow, and this soon dissipates to lower values. The drag coefficient also expresses an oscillatory behaviour, although the amplitude is small. After great fluctuations in the exploration stage, the drag coefficient converges to a stable value, which is also lower than the initial drag and expresses oscillatory behaviour yet, with reduced amplitude. Figure 11 shows the reduction in the amplitude of oscillations in the lift force. When the time-averaged mean speed is calculated in the flow field along the longitudinal center line, the reduction after control is evident and shown in Figure 12. A more important observation is the reduction in time-averaged standard deviation of flow speed along the downstream. The calculated field is a measure of averaged fluctuation shown in Figure 13. A clear evidence of how the fluctuations are minimized due to the control is captured. The fluctuation is also measured along the same longitudinal center line to get a clearer idea of the magnitude of fluctuations before and after the control, which is Figure 14.
7 Conclusion
@@
This research underscores the potential of DRL algorithms with jet actuators, and sensor arrays as an add-on in active flow control. The findings open up new avenues for optimizing flow control in practical engineering applications and hold promise for reducing noise, drag, and enhancing the performance of various engineering systems. Future work in this area can explore more complex flow scenarios, further refine control strategies, and investigate the application of DRL in other engineering domains. The study serves as a stepping stone towards the integration of machine learning techniques for enhancing the efficiency and performance of active flow control systems.

### v3

@@ III Problem introduction and setup
The uniformly separated measurements around the cylinder can be used to determine the drag coefficient by summation of discrete measurements as approximate integration. For further details about the dimensions of the setup, refer to Figure˜2.
@@
$u(y)=\frac{4Uy(0.41-y)}{0.41^{2}}$ (6)
$y$ is the y-coordinates, and $U$ is $1.5~ms^{-1}$ in this scenario, instead of the sinusoidal profile in the test problem as provided by Turek Schäfer1996; John (2004). So, the mean velocity
$\overline{U}=\frac{2}{3}U=1.0~ms^{-1}.$ (7)
Furthermore, the outflow is the rightmost wall. The upper, lower, and obstacle walls all have a non-slip condition ($u=0$) as presented in Figure˜2. The kinematic viscosity is 0.001 and diameter ($D$) of the cylinder is $0.1$. So, the Reynolds number,
$Re=\frac{\overline{U}D}{\nu}=100.$ (8)
IV Jets Configuration
Two jets with blowing and suction control are used to manipulate the flow. The first jet (referred to as Jet 1) is at the top of the circular base of the cylinder (at coordinate ($0.2$, $0.25$)), and the second jet (Jet 2) is at the bottom (at coordinate ($0.2$, $0.15$)). The width ($w$) of the jets is small, at $0.25$ percent of the diameter, i.e. 2.5e4m2.5e{-4}~m. The jets can perform blowing and suction independently, meaning blowing and suction can happen simultaneously. A reinforcement learning algorithm is applied to control the blowing and suction of the jets. More information about the execution of the simulation will be provided in the next section.
V Feedback Formulation
@@
To measure the SPLs, pressure is recorded from the pressure field provided in the simulation with surfaces of closely located sensors around the cylinder, 0.05m away from the container’s walls (refer to Figure˜3). The sensor surfaces enclose the vortex street created by the flow and therefore can give a more accurate reflection of the varying pressure field along the vortex street and it helps us determine the static pressure level, which are essential for the calculation of the SPLs. We set 2000 sensor points horizontally and 500 sensor points vertically on each side. The pressure of each point at a particular time is then extracted from the pressure field created by the simulation. The upper and lower sensor surfaces are mainly concerned as they cover the length of the vortex street, which will be the source of most of the noise generation. These vortices are born from instabilities in the bottom and top regions of flow separation alternatively. Hence, having a distinction between the top horizontal sensor array and the bottom horizontal sensor array helps in understanding the vortex periodicity. Pressure values of every sensor point at each time step are recorded and passed through a function to convert to the relative $SPL_{i}$ for each sensor and effective SPL for the system $SPL_{eff}$ using the formulae below.
$SPL_{i}=20\mathrm{log}\frac{|p_{i}-p_{avg}|_{rms}}{p_{avg}}$ (9)
$SPL_{eff}=10\mathrm{log}\frac{\sum_{i=1}^{n}|p_{i}-p_{avg}|^{2}_{rms}}{p^{2}_{avg}}$ (10)
@@ VI Experiments and Discussions
VI.1 Stage 1: Explore the optimal range of jet values
@@
VI.1.1 Set-up
It is challenging to determine what velocity value of the jet is able to influence the flow as we desired. Therefore, build-up strategy is used for the DQN to explore the value that can lower the SPLs. The velocity values of the jets are generated based on the values of the previous time step. The increments and decrements include ±0.01, ±0.05, ±0.1, ±0.5 and 0, so there shall be 81 combinations of actions for the DQN algorithm to manage. For example, if both jets initially have the values of 0.1, the algorithm will have the option to decrease Jet 1 by 0.01, 0.05, 0.1, 0.5 or keep it constant, and likewise for Jet 2. This results in a large number of actions for the DQN to explore, which can potentially hinder the ability to optimize the reward as the optimization problem needs more variables to be tuned and the gradient surface becomes rough and noisy, though the capability of optimization with more variables is better. Hence, fewer controlling variables are typically preferred. The velocity values are kept constant between the two interventions. The input of the DQN is a vector with a dimension of 1x4. The first two are for the SPLs of upper and lower sensor surfaces and the last two are the velocity values of Jet 1 and Jet 2. In this stage, we follow the model in Figure˜1, which allows the simulation to pass the data (reward, action, states) to the DQN at every time step. However, the jet velocity and therefore the action are kept constant every 50 time steps.
The reward returned is determined by a function in the simulation. In this task, the simulation learns how to reduce approximately 3 - 5 dB and create a convergence in the process. A simulation runs without any jet intervention results in two oscillations. While the lower SPL peaks at roughly 74.5 dB, the upper SPL’s peak is about 0.25 dB lower. Moreover, the range of the SPL is from 71.7 to 74.5 dB (2.8 from maximum to minimum). Refer to Figure˜5 for further information. Meanwhile, the drag coefficient is also measured due to its popularity in wake control, so this measurement can be used as a means of verification for the noise reduction method. The drag experienced initially is large due to the direction of flow at the beginning. When the flow stabilizes, the plot suggests it has an oscillatory behaviour, with a maximum and a minimum at approximately 3.185 and 3.123, respectively.
@@ VI.1.1 Set-up
VI.1.2 Results
After applying the jet intervention and running the simulation to its completion, the result is plotted in Figure˜6. From t = 4 to t = 6 on the x-axis, the SPL is stable because there is no intervention yet. However, from t = 6, there are many fluctuations in the SPL due to the exploration of the DQN algorithm. Many large values of jet values are chosen, which leads to observable discrepancies before the 8-second mark. The overall SPL in this time range is still relatively the same as when the jet is not turned on, but there is a small surge near t = 8, demonstrating the model has been able to achieve a jet pattern that can affect the SPL oscillations.
From t = 8 to t = 10, the SPL continues to rise to a peak of 75 dB, then gradually falls over time. It is also evident that there are some interventions that can bring the SPL down, preventing it from increasing drastically. The act of exploration searches for optimal policy control by letting random actions to learn the optimality like mutation steps in the genetic algorithm and simultaneously exploits the learning with the actions that reduce the SPL by preventing the random actions letting more actions from optimal policy to learn optimality. This causes the initial SPL to rise and then gradually dropKaelbling et al. (1996); Burnetas and Katehakis (1997). The SPL keeps its decreasing tendency until t = 16, as the interventions are noticeably fewer as the DQN agent slowly enters the exploiting phase.
@@ VI.1.2 Results
Looking at the velocity plot in Figure˜6b, it can be seen that a velocity with a magnitude of around 1.0 can influence the flow to the extent we desire. Along with results from other trials, one of which will be presented in Figure˜7, it can be deduced that jet velocity with a magnitude above 3 shall likely cause simulation failure due to the breakdown of the PDE solver at high jet velocities due to the instability of the discretization scheme because of the high courant number near the jets. Essentially, it is a computational limitation and it can be tackled by using higher-order numerical schemes or finer discretization. In Figure˜7a, although the SPL is at the desired value, the simulation is stopped at just past t =10 (5000 mark on the x-axis). So, the lower value of the jet is at -3.5 and that is not ideal for the simulation. This threshold will be used to limit the jet speed in Stage 2 for the test cases.
VI.2 Stage 2: Testing with definite jet velocity values
VI.2.1 Set-up
@@
Based on the result of the previous section, we set a deterministic DQN algorithm with blowing and suction of three different values. In this stage, the jet velocity is no longer built up from the previous time step but rather has a distinct, fixed set of values. Two different test cases are presented: [±1.50, ±2.25, ±2.75] and [±2.00, ±2.50, ±3.00] (referred to as Test case 1 and Test case 2 respectively), to verify the result. The expectation from this stage is similar to the one before, which is lowering from 3 - 5 dB approximately. The reward calculation function is kept the same as in the previous stage (as in Table 1), but the reward calculation process is different. While the previous stage fully adopts the Markov model (Figure˜1), the strategy for this stage is accumulating the rewards between two interjections and averaging them, then returning to the DQN. Lastly, the 50-divisible state is returned to the DQN. Both test cases will implement this model. Refer to Figure˜8 for further details.
VI.2.2 Results
The resulting SPL of the upper and lower sensors are displayed in Figure˜9 for both cases. Test case 1, using velocity values with lower magnitudes, converges to an SPL of just below 70 dB. Meanwhile, Test case 2 demonstrates a slightly lower SPL than that of Test case 1, and it seems to converge earlier as well. Also, the amplitude of oscillations is evident to have reduced.
On the other hand, the instantaneous drag coefficient in each case is calculated and the result is impressive(refer to Figure˜10). The system experiences a large drag force at the start due to the direction of the flow, and this soon dissipates to lower values. The drag coefficient also expresses an oscillatory behaviour, although the amplitude is small. After great fluctuations in the exploration stage, the drag coefficient converges to a stable value, which is also lower than the initial drag and expresses oscillatory behaviour yet, with reduced amplitude. Figure˜11 shows the reduction in the amplitude of oscillations in the lift force. When the time-averaged mean speed is calculated in the flow field along the longitudinal center line, the reduction after control is evident and shown in Figure˜12. A more important observation is the reduction in time-averaged standard deviation of flow speed along the downstream. The calculated field is a measure of averaged fluctuation shown in Figure˜13. A clear evidence of how the fluctuations are minimized due to the control is captured. The fluctuation is also measured along the same longitudinal center line to get a clearer idea of the magnitude of fluctuations before and after the control, which is Figure˜14.
@@ VI.2.2 Results
VI.3 Validation
VI.3.1 Computational Framework
Validation and numerical sensitivity analysis are necessary to establish the reliability of the computational framework. To do so, specifically, we conducted mesh independence analysis (Table˜2) and time-step sensitivity analysis (Table˜3).
Multiple mesh densities and time-step sizes were tested, and the resulting drag coefficient, and lift coefficient were monitored. The selected numerical parameters correspond to the regime where variations in the key flow quantities became negligibly small.
Mesh Elements $C_{D}$ $C_{L}$ (rms)
Coarse 42,000 3.08 0.212
Medium 85,000 3.14 0.226
Fine 168,000 3.15 0.229
Table 2: Mesh independence study for the uncontrolled cylinder flow case.
$\Delta t$ $C_{D}$ $C_{L}$ (rms) Relative Error (%)
0.005 3.11 0.221 1.27
0.002 3.14 0.226 0.31
0.001 3.15 0.226
Table 3: Time-step sensitivity analysis for the uncontrolled flow case.
Case $C_{D}$ (mean) Red. (%) $C_{D}$ (std) Red. (%) $C_{L}$ (std) Red. (%) SPL (dB) Red. (%)
Baseline (No Control) 3.15 0 0.031 0 0.23 0 73.5 0
DPM Adjoint (L1H100, Liu and MacArt (2024)) 2.35 25.4 0.0003 99.0 0.006 97.4 - -
DRL using Velocity (L1H100, Liu and MacArt (2024)) 3.20 - 1.6 0.024 22.6 - - - -
DRL using Velocity (L2H512, Liu and MacArt (2024)) 3.03 3.8 0.014 54.8 - - - -
DRL using SPL (ours, test case 1) 2.65 15.9 0.018 41.9 0.15 34.8 69.0 6.1
DRL using SPL (ours, test case 2) 2.40 23.8 0.012 61.3 0.11 52.2 66.5 9.5
Table 4: Acoustic-based DRL-based flow control against baseline and benchmark data for flow past a circular cylinder.
VI.3.2 Benchmarking
Table˜4 further demonstrates that the acoustically informed DRL framework substantially outperforms velocity-based DRL configurations in terms of drag reduction and fluctuation suppression. This observation suggests that hydrodynamic pressure feedback provides a more physically informative control signal for wake stabilization under the present flow conditions.
VI.3.3 Statistical Convergence
Time averaging was performed only after the control policy reached a statistically stationary regime following the initial exploration phase. To verify the statistical convergence of the time-averaged flow quantities, convergence analyses of the mean velocity and vorticity fields were performed using progressively increasing averaging windows after the transient regime had decayed. The time-averaged quantity was computed as:
$\overline{\phi}=\frac{1}{T}\int_{t_{0}}^{t_{0}+T}\phi(t)\,dt$ (11)
where $\phi$ denotes either the velocity or vorticity field, and $T$ represents the averaging duration. To assess statistical convergence, the relative variation between successive averaging windows was evaluated as:
$\epsilon(T)=\frac{\left\|\overline{\phi}_{T+\Delta T}-\overline{\phi}_{T}\right\|}{\left\|\overline{\phi}_{T}\right\|}$ (12)
where $\|\cdot\|$ denotes the appropriate spatial norm. It was observed that both the mean velocity and vorticity fields reached statistically converged states with relative variations below $1\%$ after approximately 20 vortex shedding cycles. Furthermore, the uncertainty associated with the averaged quantities was quantified using the temporal standard deviation:
$\sigma_{\phi}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\phi_{i}-\overline{\phi})^{2}}$ (13)
and the standard error of the mean (SEM):
$\mathrm{SEM}=\frac{\sigma_{\phi}}{\sqrt{N}}$ (14)
where $N$ denotes the total number of statistically independent samples. The resulting uncertainty levels were found to be sufficiently small compared to the corresponding mean values, confirming the robustness and statistical reliability of the reported flow quantities. See the results in Table˜5.
Quantity Mean Value Std. Dev. SEM
Streamwise Velocity $\overline{u}$ 0.842 0.021 0.0021
Vorticity Magnitude $\overline{\omega}$ 5.37 0.18 0.018
Table 5: Statistical convergence and uncertainty analysis of time-averaged flow quantities.
VI.3.4 Sensor Placement
Since the proposed control framework relies on pressure-based acoustic feedback, the spatial distribution and density of the sensor measurements can indeed influence the observed control behavior.
In the present study, the sensor surfaces were intentionally positioned around the wake region in order to capture the dominant hydrodynamic pressure fluctuations associated with vortex shedding. The upper and lower sensor arrays were selected specifically to monitor the alternating pressure signatures generated by the von Kármán vortex street.
The relatively dense sensor distribution was adopted to obtain a more accurate spatially averaged estimate of the wake-induced pressure fluctuations and reduce sensitivity to localized numerical oscillations.
To assess the robustness of the framework, additional tests were performed using reduced sensor densities and modified sensor locations. The overall control trends, including suppression of wake fluctuations and drag reduction, remained qualitatively consistent, although moderate quantitative variations in SPL reduction and convergence behavior were observed. See Table˜6.
These observations suggest that while sensor placement influences the sensitivity and responsiveness of the control policy, the proposed framework is not exclusively dependent on a single sensor configuration.
Sensor SPL Drag Convergence
Configuration Reduction (%) Reduction (%) Stability
Coarse
(500 $\times$ 100) 7.8 18.5 Moderate
Medium
[…]
Fine
(2000 $\times$ 500) 9.5 23.8 Good
Table 6: Sensitivity of control performance to sensor density.
VII Conclusion
The results demonstrate that acoustically informed DRL-guided synthetic jet actuation can effectively attenuate wake fluctuations, reduce hydrodynamic pressure oscillations, and suppress drag-generating vortex structures in laminar cylinder wake flow. The framework further illustrates the feasibility of coupling reinforcement learning with pressure-based sensing for adaptive active flow control under nonlinear wake dynamics.
@@ VII Conclusion
Although the present framework does not seek globally optimal control in the strict optimal-control sense, the DRL-guided strategy consistently identifies dynamically effective actuation regimes capable of reducing wake fluctuations, aerodynamic drag, and acoustic pressure levels. The results demonstrate the feasibility of acoustically informed reinforcement learning for adaptive wake stabilization under nonlinear flow conditions. This research underscores the potential of DRL algorithms with jet actuators, and sensor arrays as an add-on in active flow control. The findings open up new avenues for optimizing flow control in practical engineering applications and hold promise for reducing noise, drag, and enhancing the performance of various engineering systems. Future work in this area can explore more complex flow scenarios, further refine control strategies, and investigate the application of DRL in other engineering domains. The study serves as a stepping stone towards the integration of machine learning techniques for enhancing the efficiency and performance of active flow control systems.